### v1

# FISH-Tuning: Enhancing PEFT Methods with Fisher Information

Abstract
The rapid growth in the parameter size of Large Language Models (LLMs) has led to the development of Parameter-Efficient Fine-Tuning (PEFT) methods to alleviate the computational costs of fine-tuning. Among these, Fisher Induced Sparse uncHanging (FISH) Mask is a selection-based PEFT technique that identifies a subset of pre-trained parameters for fine-tuning based on approximate Fisher information. However, the integration of FISH Mask with other PEFT methods, such as LoRA and Adapters, remains underexplored. In this paper, we propose FISH-Tuning, a novel approach that incorporates FISH Mask into addition-based and reparameterization-based PEFT methods, including LoRA, Adapters, and their variants. By leveraging Fisher information to select critical parameters within these methods, FISH-Tuning achieves superior performance without additional memory overhead or inference latency. Experimental results across various datasets and pre-trained models demonstrate that FISH-Tuning consistently outperforms the vanilla PEFT methods with the same proportion of trainable parameters.  [^1] The code for this work will be made openly accessible after the anonymous review process.
FISH-Tuning: Enhancing PEFT Methods with Fisher Information
Kang Xue1,2,3, Ming Dong1,2,3, Xinhui Tu1,2,3, Tingting He1,2,3*, 1Hubei Provincial Key Laboratory of Artificial Intelligence and Smart Learning 2National Language Resources Monitoring and Research Center for Network Media 3School of Computer, Central China Normal University, Wuhan, China *Corresponding author xuekang@mails.ccnu.edu.cn {dongming,tuxinhui,tthe}@ccnu.edu.cn
1 Introduction
The emergence of Large Language Models (LLMs) has revolutionized Natural Language Processing (NLP) by achieving remarkable performance across a wide range of tasks. These models, typically trained on massive datasets using self-supervised learning, are fine-tuned on downstream tasks through processes such as Supervised Fine-Tuning (SFT) and Reinforcement Learning with Human Feedback (RLHF) Christiano et al. (2017); Stiennon et al. (2020); Ouyang et al. (2022). However, fine-tuning all parameters of such large LLMs is computationally expensive, requiring substantial GPU memory and training time. This challenge has led to the rise of Parameter-Efficient Fine-Tuning (PEFT) Ding et al. (2023) methods, which aim to achieve competitive performance by fine-tuning only a small subset of parameters.
PEFT methods can be broadly categorized into three types Lialin et al. (2023); Han et al. (2024): (1) Selection-based methods, which fine-tune a subset of pre-trained parameters while freezing the rest, e.g., BitFit Zaken et al. (2022), Diff-Pruning Guo et al. (2021), and FISH Mask Sung et al. (2021). (2) Addition-based methods, which introduce additional trainable parameters or layers into the model, e.g., Adapters Houlsby et al. (2019), $(\text{IA})^{3}$ Liu et al. (2022), and Prefix-Tuning Li and Liang (2021). (3) Reparameterization-based methods, which use low-rank representations to reduce the number of trainable parameters, e.g., LoRA Hu et al. (2022), DoRA Liu et al. (2024), and IntrinsicSAID Aghajanyan et al. (2021). While these methods have proven effective individually, hybrid approaches that combine multiple PEFT techniques are gaining attention to further improve efficiency and performance.
FISH Mask, a selection-based PEFT method, identifies the most critical parameters for fine-tuning using Fisher information. Despite its promise, its integration with addition-based and reparameterization-based methods remains underexplored. For instance, LoRA introduces trainable low-rank weight matrices while freezing the pre-trained model weights. These matrices could also benefit from Fisher information to identify the most important parameters, potentially enhancing LoRA’s performance.
In this paper, we propose FISH-Tuning, a novel framework that integrates FISH Mask into addition-based and reparameterization-based PEFT methods, including LoRA, Adapters, and their variants. FISH-Tuning leverages Fisher information to select the most critical parameters within these methods, enabling efficient fine-tuning without additional memory overhead or inference latency. Specifically, we demonstrate how FISH Mask can be applied to LoRA, DoRA, Adapters, Prefix-Tuning, and $(\text{IA})^{3}$. Experimental results across multiple datasets and pre-trained models show that FISH-Tuning consistently outperforms the original PEFT methods with the same proportion of trainable parameters. We summarize our contributions as follows:
  • We introduce FISH-Tuning, a novel framework that integrates FISH Mask into addition-based and reparameterization-based PEFT methods, enabling efficient parameter selection without increasing memory or latency.
  • We demonstrate the effectiveness of FISH-Tuning across various datasets and pre-trained models, achieving consistent performance improvements over the original PEFT methods.
  • We provide insights into the role of Fisher information in parameter selection, offering a new perspective on optimizing PEFT methods.
2 Related Work
2.1 Parameter Efficient Fine-tuning
With the growing size of pre-trained Large Language Models, fine-tuning all parameters is becoming increasingly expensive and may lead to overfitting Mahabadi et al. (2021). Parameter-Efficient Fine-Tuning (PEFT) has been proposed to alleviate this issue. It trains only a small proportion of the parameters while achieving similar results compared to full fine-tuning.
In Transformer architecture Vaswani et al. (2017), the Multi-Head Attention mechanism is defined as:
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W^{O}$
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},KW_{i}^{K},VW_{i}^{V})$ (2)
The feed-forward network (FFN) is given by:
$\text{FFN}(\boldsymbol{x})=\max(0,\boldsymbol{x}W_{1}+b_{1})W_{2}+b_{2}$ (3)
where $Q\in\mathbb{R}^{d},K\in\mathbb{R}^{d},V\in\mathbb{R}^{d},W_{i}^{Q}\in\mathbb{R}^{d\times k},W_{i}^{K}\in\mathbb{R}^{d\times k},W_{i}^{V}\in\mathbb{R}^{d\times k},W^{O}\in\mathbb{R}^{hk\times d},\boldsymbol{x}\in\mathbb{R}^{d},W_{1}\in\mathbb{R}^{d\times r},W_{2}\in\mathbb{R}^{r\times d},b_{1}\in\mathbb{R}^{r},b_{2}\in\mathbb{R}^{d}$. Different PEFT methods modify different weights in the Transformer. In LoRA and DoRA, $W_{Q}$, $W_{K}$, $W_{V}$, $W_{O}$, $W_{1}$, and $W_{2}$ can be used for matrix Reparameterization. The Adapter method adds an additional feed-forward layer to the Transformer. Prefix-Tuning introduces prefix matrices before $KW_{i}^{K}$ and $VW_{i}^{V}$. $(\text{IA})^{3}$ incorporates trainable vectors in $KW_{i}^{K}$, $VW_{i}^{V}$, and $\max(0,xW_{1}+b_{1})$. BitFit selects only the bias terms as trainable parameters.
2.2 Fisher Information
The Fisher Information Matrix (FIM) Fisher (1922); Amari (1996) is a fundamental tool in deep learning neural networks that measures parameter importance and helps address catastrophic forgetting French (1999); McCloskey and Cohen (1989); McClelland et al. (1995); Ratcliff (1990). Its three key advantages are Pascanu and Bengio (2014): approximates the Hessian matrix near loss function minima; can be computed efficiently using first-order derivatives; is positive semi-definite, ensuring stable optimization. These properties make it particularly valuable for techniques like Elastic Weight Consolidation (EWC), which uses FIM to identify and protect important parameters while learning new tasks, thus helping preserve previously learned information.
3 Problem Statement
3.1 Task Definition
The goal of PEFT is to fine-tune the model with as few trainable parameters as possible while achieving better results. Therefore, we can compare our method with the original PEFT method using the same trainable parameter ratio and evaluate their performance on the same dataset and hyperparameters.
3.2 FISH Mask based PEFT
Fisher Information Matrix (FIM) is defined as:
$F_{\theta}=\mathbb{E}_{x\sim p(x)}[\mathbb{E}_{y\sim p_{\theta}(y|x)}\nabla_{\theta}\log p_{\theta}(y|x)\nabla_{\theta}\log p_{\theta}(y|x)^{\textrm{T}}]$
where $x$ is the input, $y$ is the output, $\theta$ represents the model’s parameters, $p(x)$ is the probability distribution of the input $x$, and $\nabla$ is the gradient. Fisher information is typically estimated using a diagonal approximation, in which gradients for all parameters are calculated based on $N$ data samples:
$\hat{F}_{\theta}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{y\sim p_{\theta}(y|x_{i})}\Big{[}\nabla_{\theta}\log p_{\theta}(y|x_{i})$
$\qquad\odot\nabla_{\theta}\log p_{\theta}(y|x_{i})\Big{]}$ (5)
where $N$ is the number of data samples, and $\odot$ is the Hadamard product. In supervised learning, we can use “Empirical Fisher information” for further approximation:
$\hat{F}_{\theta}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})\odot\nabla_{\theta}\log p_{\theta}(y_{i}|x_{i})$
We will select the top-k parameters $\theta_{i}$ according to the estimated FIM:
$\theta_{\text{selected}}=\{\theta_{i}\ |\ \hat{F}_{\theta_{i}}\geq\text{sort}(\hat{F}_{\theta})_{k}\}$ (7)
Then we create the binary mask $M$ based on the top-k importance values in $\hat{F}_{\theta}$:
$M_{i}=\begin{cases}1,&\text{if }\theta_{i}\in\theta_{\text{selected}}\\ 0,&\text{otherwise}\end{cases}$ (8)
Finally, we mask the gradients for the loss function:
$\nabla_{\theta_{i}}\mathcal{L}^{\text{masked}}=(\nabla_{\theta_{i}}\mathcal{L})\odot M_{i}$ (9)
The masked gradients can be used to update the parameters $\theta_{i}$ using Stochastic Gradient Descent (SGD) or Adam Kingma and Ba (2015) optimizer.
4 Method
We use FISH Mask into the Addition-based methods like Adapter, Prefix-Tuning, $(\text{IA})^{3}$, and the Reparameterization-based methods like LoRA, DoRA. We also use FISH Mask into the Hybrid PEFT method like UniPELT Mao et al. (2022). We believe that our method can also be used in other PEFT methods.
4.1 FISH Mask in Reparameterization-based methods
4.1.1 FISH Mask in LoRA
Figure 1: Original LoRA method (left) and the LoRA-FISH method (right).
In the original LoRA method, the update to the weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ is represented as:
$W_{0}+\Delta W=W_{0}+BA$ (10)
where $B\in\mathbb{R}^{d\times r}$ and A $\in\mathbb{R}^{r\times k}$ , and the update involves training the matrices $B$ and $A$ while $W_{0}$ is frozen.
For the FISH-Tuning method in LoRA, we define the combined vector of B and A as $\tilde{\theta}\in\mathbb{R}^{d\times r+r\times k}$. Then we use Eq. 6 to calculate the importance score of $\tilde{\theta}$ and create the related binary mask for it.
The difference between the original LoRA method and the FISH-Tuning method is shown in Fig. 1.
4.1.2 FISH Mask in DoRA
In the original DoRA method, the update to the weight matrix $W_{0}\in\mathbb{R}^{d\times k}$ is represented as:
$W_{0}+\Delta W=m\frac{W_{0}+BA}{\left\|W_{0}+BA\right\|_{c}}$ (11)
where $m\in\mathbb{R}^{1\times k}$, $B\in\mathbb{R}^{d\times r}$ , $A\in\mathbb{R}^{r\times k}$, .$\left\|.\right\|_{c}$ is the vector-wise norm of a matrix across each column , and the update involves training the matrices $B$, $A$ and vector $m$ while $W_{0}$ is frozen.
For the FISH-Tuning method in DoRA, we define the combined vector of $B$, $A$, and $m$ as $\tilde{\theta}\in\mathbb{R}^{1\times k+d\times r+r\times k}$. Then we use Eq. 6 to calculate the importance score of $\tilde{\theta}$ and create the related binary mask for it.
4.2 FISH Mask in Addition-based methods
4.2.1 FISH Mask in Adapter
Figure 2: Original Adapter method (left) and the Adapter-FISH method (right).
In the original Serial Adapter method, it adds the adapter module twice to each Transformer layer: after the projection following multi-head attention and after the two feed-forward layers. The formula can be represented as:
$\text{Adapter}(\boldsymbol{x})=B\sigma(A\boldsymbol{x})+\boldsymbol{x}$ (12)
where $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times d}$, $\boldsymbol{x}\in\mathbb{R}^{d}$, $\sigma$ is the non-linear activation function. The update involves training the matrices $B$ and $A$ while $x$ is frozen.
For the FISH-Tuning method in Adapter, we define the combined vector of $B$ and $A$ as $\tilde{\theta}\in\mathbb{R}^{d\times r+r\times d}$. Then we use Eq. 6 to calculate the importance score of $\tilde{\theta}$ and create the related binary mask for it.
The difference between the original Adapter method and the FISH-Tuning method is shown in Fig. 2.
4.2.2 FISH Mask in Prefix-Tuning
Prefix Tuning introduces new parameters into the multi-head attention blocks in each Transformer layer. More specifically, it prepends trainable prefix vectors $P^{K}$ and $P^{V}$ to the keys and values of the attention head input, each with a configurable prefix length $l$:
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},[P_{i}^{K},KW_{i}^{K}],[P_{i}^{V},VW_{i}^{V}])$
where $W_{i}\in\mathbb{R}^{d\times k}$, $Q,K,V\in\mathbb{R}^{seq\_len\times d}$, $P_{i}\in\mathbb{R}^{l\times k}$, and $[P_{i}^{K},KW_{i}^{K}]\in\mathbb{R}^{(l+seq\_len)\times k}$. $head_{i}$ is the $i$-th attention head. The update involves training the matrices $P_{i}$ while $W_{i}$ is frozen.
For the FISH-Tuning method in Prefix-Tuning, we define the combined vector of $P_{i}^{K}$ and $P_{i}^{V}$ as $\tilde{\theta}\in\mathbb{R}^{2\times l\times k}$. Then we use Eq. 6 to calculate the importance score of $\tilde{\theta}$ and create the related binary mask for it.
4.2.3 FISH Mask in $(\text{IA})^{3}$
In the original $(\text{IA})^{3}$ method, it introduces trainable vectors into different components of a Transformer model, which perform element-wise rescaling of inner model activations. The formula can be represented as:
$\text{head}_{i}=\text{Attention}(QW_{i}^{Q},l_{k}\odot KW_{i}^{K},l_{v}\odot VW_{i}^{V})$
$\text{FFN}(x)=(l_{ff}\odot\sigma(Ax))B$ (15)
where $W_{i}\in\mathbb{R}^{d\times k}$, $Q,K,V\in\mathbb{R}^{seq\_len\times d}$, $B\in\mathbb{R}^{d\times r}$, $A\in\mathbb{R}^{r\times d}$, $x\in\mathbb{R}^{d\times k}$, $l\in\mathbb{R}^{k}$, and $l_{k}\odot KW_{i}^{K}\in\mathbb{R}^{seq\_len\times k}$. $head_{i}$ is the $i$-th attention head. $FFN$ is the Feed-forward Network. $\sigma$ is the non-linear activation function. The update involves training the vectors $l$ while $W_{i}$, $A$, and $B$ are frozen.
For the FISH-Tuning method in $(\text{IA})^{3}$, we define the combined vector of $l_{k}$, $l_{v}$ and $l_{ff}$ as $\tilde{\theta}\in\mathbb{R}^{3\times k}$. Then we use Eq. 6 to calculate the importance score of $\tilde{\theta}$ and create the related binary mask for it.
4.3 FISH Mask in UniPELT
In the original UniPELT method, it adds a trainable gating value ${\mathcal{G}_{m}}\in(0,1)$ that is computed via a feed-forward network $W_{\mathcal{G}_{m}}$ and sigmoid activation $\sigma$ from the Transformer layer input states $x$:
$\mathcal{G}_{m}=\sigma(W_{\mathcal{G}_{m}}x)$ (16)
These gating values are then used to scale the output activations of the injected PEFT modules, e.g., for a LoRA layer:
$W_{0}+\Delta W=W_{0}+\mathcal{G}_{LoRA}BA$ (17)
where the update involves training the matrices $W_{\mathcal{G}_{m}}$, $B$, and $A$ while $W_{0}$ is frozen.
In our settings for UniPELT, we use LoRA, Adapter, Prefix-Tuning as modules and add separate gating values for them. For the FISH-Tuning method, we follow the settings from Character 4.1.1 for the LoRA component, Character 4.2.1 for the Adapter component, and Character 4.2.2 for the Prefix-Tuning component.
5 Experiments Setup
5.1 Datasets and Baselines
Datasets. We evaluate FISH-Tuning method on the GLUE Wang et al. (2019) dataset, comparing it with the original PEFT method mentioned in Chapter 4. GLUE is a multi-task benchmark that contains 10 datasets for LLM evaluation. In our experiment, we select only the CoLA Warstadt et al. (2018), MRPC Dolan and Brockett (2005), RTE, SST-2 Socher et al. (2013), STS-B Cer et al. (2017), and WNLI Levesque et al. (2012) datasets because the remaining datasets contain too much text and require excessive training time. Then we calculate the average score of these six datasets.
Method Trainable Parameters CoLA MRPC RTE SST-2 STS-B WNLI Avg
Original-LoRA 0.0057% 43.27 80.94 58.48 89.56 84.94 53.52 68.45
LoRA-FISH 0.0057% 44.28 80.25 58.48 90.48 86.41 53.52 68.90
[…]
LoRA-FISH 0.0099% 51.21 85.74 66.06 90.14 86.84 53.52 72.25
[…]
LoRA-FISH 0.0142% 53.58 85.56 65.70 90.71 86.79 53.52 72.64
[…]
LoRA-FISH 0.0184% 55.96 84.18 67.51 89.45 86.81 53.52 72.91
Original-DoRA 0.0078% 42.70 80.45 58.48 89.68 85.04 53.52 68.31
DoRA-FISH 0.0078% 46.36 80.30 58.48 91.17 86.84 53.52 69.45
[…]
DoRA-FISH 0.0142% 54.14 87.47 66.43 90.14 86.84 53.52 73.09
[…]
DoRA-FISH 0.0206% 55.89 85.63 66.79 89.91 86.86 53.52 73.10
[…]
DoRA-FISH 0.0269% 56.00 86.70 67.51 89.91 87.21 53.52 73.48
Original-Adapter 0.1389% 41.53 78.54 62.82 85.32 82.54 56.34 67.85
Adapter-FISH 0.1389% 52.27 87.36 64.26 90.60 87.26 57.75 73.25
[…]
Adapter-FISH 0.2760% 53.73 88.91 64.62 90.48 87.62 57.75 73.85
[…]
Adapter-FISH 0.4127% 52.64 87.76 65.70 91.28 87.76 57.75 73.82
[…]
Adapter-FISH 0.5490% 53.13 88.19 64.26 91.51 87.98 57.75 73.80
Original-PrefixTuning 0.0439% 34.03 76.57 59.57 83.83 80.65 57.75 65.40
PrefixTuning-FISH 0.0439% 38.17 76.60 60.29 88.07 82.92 59.15 67.54
[…]
PrefixTuning-FISH 0.0864% 39.22 77.15 60.29 87.61 83.94 59.15 67.90
[…]
PrefixTuning-FISH 0.1289% 38.91 77.13 60.29 87.61 84.47 59.15 67.93
[…]
PrefixTuning-FISH 0.1713% 40.94 77.13 60.29 88.76 84.69 59.15 68.49
Original-$(\text{IA})^{3}$ 0.0142% 34.11 76.42 64.62 87.16 84.68 47.89 65.81
$(\text{IA})^{3}$-FISH 0.0142% 38.54 78.03 64.62 89.11 86.60 47.89 67.46
[…]
$(\text{IA})^{3}$-FISH 0.0227% 38.47 81.28 64.26 88.99 87.04 47.89 67.99
[…]
$(\text{IA})^{3}$-FISH 0.0354% 37.99 80.06 62.09 89.11 87.41 46.48 67.19
[…]
$(\text{IA})^{3}$-FISH 0.0439% 46.74 81.26 62.45 89.22 87.59 46.48 68.96
Original-UniPELT 0.0213% 47.25 81.76 60.65 90.71 85.65 66.20 72.04
UniPELT-FISH 0.0213% 49.72 81.28 64.26 90.14 86.94 43.66 69.33
[…]
UniPELT-FISH 0.0411% 55.70 81.95 63.90 91.74 87.39 43.66 70.72
[…]
UniPELT-FISH 0.0610% 51.66 87.38 63.90 91.28 87.45 43.66 70.89
[…]
UniPELT-FISH 0.0808% 52.09 87.80 63.90 90.60 87.45 43.66 70.92
Table 1: Performance of different methods on different datasets. The solid lines separate different PEFT methods, while the dashed lines separate different ratios of trainable parameters. In each dashed-line area, the first row represents the original method, and the second row represents our method.
Due to the limitation of uploading test set results to the official website only twice a day, we use only the validation set results.
Baselines. We compare FISH-Tuning method with the original PEFT method using the same dataset and hyperparameters. We use various datasets, PEFT methods, and pre-trained models to demonstrate that FISH-Tuning is better than the original one.
5.2 Evaluation Metrics
Different datasets have different evaluation metrics. The CoLA dataset uses Matthews correlation coefficient Matthews (1975). The STS-B uses Pearson and Spearman correlation coefficients. The MRPC uses a combined score (half the sum of $F_{1}$ and Accuracy). The rest of the datasets use Accuracy. We also analyze the average loss score of these six datasets.
5.3 Implementation Details
We follow the same parameter setting for the classification task as BERT Devlin et al. (2019). The initialization weight method of the matrices $B$ and $A$ in LoRA and DoRA follows PiSSA Meng et al. (2024). For experiments, we did not use any hyper-parameter tuning, nor did we use MLNI trick (use the MLNI checkpoint instead of the pre-trained weights) to enhance the models’ performance. More details about the hyperparameters are available in Table 6 in Appendix C.
For different trainable parameter ratios, we select various layers of the LLM as trainable parameters to achieve different trainable ratios. For most PEFT methods, we select the bottom 1, 2, 3, and 4 layers to form four groups with different trainable ratios. In FISH-Tuning method, we select the bottom 5 layers as the parameter set and use the FISH Mask to choose the top-k parameters from this set, ensuring the same trainable ratios as the original PEFT method. For $(\text{IA})^{3}$, we select the bottom 3, 5, 8, and 10 layers to form four groups with different trainable ratios, while in the FISH-Tuning method, we use 12 layers because $(\text{IA})^{3}$ trains fewer parameters compared to other PEFT methods.
5.4 Experimental Results
5.4.1 Baselines
We conduct extensive experiments on the GLUE dataset and the BERT model to verify the effectiveness of FISH-Tuning. Table 1 presents the detailed performance of FISH-Tuning and baseline methods on the GLUE benchmark. Overall, the results demonstrate that integrating the FISH Mask into various PEFT methods consistently improves performance across multiple tasks, except for UniPELT.
In LoRA, FISH-Tuning consistently outperforms the original method. At a low parameter ratio of 0.0057%, while the improvement on MRPC and RTE is minimal, FISH-Tuning achieves higher scores on CoLA, SST-2, and STS-B, resulting in an overall average boost (from 68.45 to 68.90). As the trainable parameter ratio increases (e.g., to 0.0099% and 0.0184%), the improvements become even more significant, with average scores rising from 70.42 to 72.25 and from 72.11 to 72.91.
In DoRA, FISH-Tuning enhances performance on nearly all tasks, with average scores increasing by up to 2.75 points at a parameter ratio of 0.0142%. Notably, significant improvements on tasks like CoLA and RTE highlight the robustness of FISH-Tuning across different parameter scales.
In Adapter, at a 0.1389% trainable parameter ratio, the original Adapter method achieves an average score of 67.85, while FISH-Tuning raises the average to 73.25—a gain of over 5 points. This trend continues at higher parameter ratios (0.2760%, 0.4127%, and 0.5490%), confirming that FISH-Tuning not only enhances performance but also scales well with an increased trainable parameter ratio.
In Prefix-Tuning, integrating the FISH Mask generally leads to notable improvements at lower trainable parameter ratios. For example, at a 0.0439% parameter ratio, FISH-Tuning raises the CoLA score from 34.03 to 38.17, SST-2 from 83.83 to 88.07, and STS-B from 80.65 to 82.92—resulting in an overall average increase from 65.40 to 67.54. Similarly, at 0.0864% and 0.1289%, the average performance improves from 65.83 to 67.90 and from 66.13 to 67.93, respectively. However, at a higher parameter ratio of 0.1713%, although gains are observed on CoLA and SST-2 (e.g., CoLA increases from 35.80 to 40.94), notable declines on tasks such as RTE (from 65.34 down to 60.29) and WNLI (from 64.79 down to 59.15) lead to an overall average drop from 69.34 to 68.49. These findings suggest that while FISH integration in Prefix-Tuning is effective at lower parameter scales, its impact becomes less consistent as more parameters are tuned.
In $(\text{IA})^{3}$, FISH-Tuning yields mixed outcomes that appear to depend on the parameter ratio. At a low ratio of 0.0142%, FISH-Tuning consistently improves performance—boosting CoLA from 34.11 to 38.54, SST-2 from 87.16 to 89.11, and STS-B from 84.68 to 86.60, which increases the average score from 65.81 to 67.46. When the ratio is increased to 0.0227%, improvements in MRPC and STS-B are offset by slight drops in CoLA and RTE, resulting in a minimal average gain (67.91 to 67.99). At 0.0354%, decreases in CoLA and RTE cause the average to drop from 67.54 to 67.19. Notably, at 0.0439%, the integration of FISH Mask leads to a significant boost in CoLA (from 39.59 to 46.74) along with modest gains in RTE and SST-2, raising the average from 67.64 to 68.96. Overall, while FISH-Tuning can enhance $(\text{IA})^{3}$ performance under certain settings, its benefits are somewhat sensitive to the trainable parameter ratio.
In contrast, the integration of the FISH Mask in UniPELT does not yield consistent benefits. At a 0.0213% trainable parameter ratio, despite improvements in CoLA (increasing from 47.25 to 49.72) and RTE (from 60.65 to 64.26), there is a dramatic drop in WNLI—from 66.20 to 43.66—which pulls the average score down from 72.04 to 69.33. This adverse trend persists across higher parameter ratios: at 0.0411%, the average falls from 72.73 to 70.72; at 0.0610%, from 73.85 to 70.89; and at 0.0808%, from 72.72 to 70.92. We give our possible hypothesis for this experiment result in Limitations.
Method Trainable Parameters CoLA MRPC RTE SST-2 STS-B WNLI Avg
Original-LoRA (M) 0.0033% 33.93 79.14 57.04 83.37 79.24 46.48 63.20
LoRA-FISH 0.0033% 36.60 80.14 57.76 86.24 81.52 53.52 65.96
[…]
LoRA-FISH 0.0056% 35.25 79.44 58.12 86.24 81.95 53.52 65.75
[…]
LoRA-FISH 0.0080% 36.44 79.97 58.48 86.58 82.66 53.52 66.28
[…]
LoRA-FISH 0.0103% 40.04 79.10 58.48 87.16 82.58 53.52 66.81
Original-LoRA (L) 0.0007% 46.80 77.31 68.59 89.68 80.71 57.75 70.14
LoRA-FISH 0.0007% 47.07 79.67 70.04 90.94 82.70 57.75 71.36
[…]
LoRA-FISH 0.0010% 52.67 80.24 72.20 91.17 83.93 56.34 72.76
[…]
LoRA-FISH 0.0013% 52.16 80.43 72.20 91.86 84.77 56.34 72.96
[…]
LoRA-FISH 0.0017% 53.84 80.45 72.56 90.71 84.29 56.34 73.03
Table 2: Performance of different tasks in GLUE. (M) means Modern BERT. (L) means LLaMA-3.2-1B.
5.4.2 Loss on Evaluation Datasets
We conduct an experiment on the loss scores using the evaluation dataset. The loss values are taken from the first solid-line area in Table 1. We visualize the loss value curves for the original PEFT method and FISH-Tuning method in Fig. 3. In this figure, the x-axis represents the training epochs, while the y-axis denotes the average evaluation loss scores across six datasets. From this figure, we observe that all points on the FISH-Tuning curve are lower than those on the original PEFT method curve. This result demonstrates that the FISH-Tuning method converges faster than the original PEFT method. It also explains why FISH-Tuning method achieves better results in Table 1.
Figure 3: Original LoRA method loss (Blue line) and the LoRA-FISH method loss (Orange line).
5.4.3 Results with other Pre-trained Models
Apart from the BERT model, we also compare FISH-Tuning with the original PEFT method using other trending pre-trained models, such as ModernBERT Warner et al. (2024) and LLaMA-3.2-1B Dubey et al. (2024). The prompt template for LLaMA-3.2-1B follows Table 1 in Zhong et al. (2023). Apart from the prompt setting, the rest of the settings are exactly the same as the BERT settings. The experiment results can be seen in Table 2.
In the ModernBERT experiments, we observe that integrating the FISH Mask into LoRA consistently improves performance across nearly all tasks. For instance, at a very low trainable parameter ratio of 0.0033%, FISH-Tuning boosts the CoLA score from 33.93 to 36.60 and enhances SST-2 from 83.37 to 86.24, leading to an overall average increase from 63.20 to 65.96. Similar trends are evident at slightly higher parameter ratios—at 0.0056% and 0.0080%, modest improvements in metrics such as RTE and MRPC are complemented by more pronounced gains in SST-2 and STS-B, yielding average scores rising from 63.50 to 65.75 and from 63.97 to 66.28, respectively. Even at 0.0103%, where the original LoRA attains an average of 66.17, FISH-Tuning further pushes the SST-2 and STS-B scores (up to 87.16 and 82.58, respectively), resulting in an average improvement to 66.81.
For LLaMA-3.2-1B, the benefits of FISH-Tuning are even more pronounced. At the lowest parameter ratio of 0.0007%, FISH-Tuning not only raises the CoLA score from 46.80 to 47.07 but also significantly improves MRPC and RTE (from 77.31 to 79.67 and 68.59 to 70.04, respectively), which lifts the overall average from 70.14 to 71.36. At a parameter ratio of 0.0010%, despite the strong baseline performance (average of 71.36), FISH-Tuning further enhances the performance to 72.76 by providing substantial gains on CoLA and RTE. With increasing parameter ratios of 0.0013% and 0.0017%, FISH-Tuning continues its steady performance improvements, achieving average scores of 72.96 and 73.03, respectively.
These results on ModernBERT and LLaMA-3.2-1B highlight the generalizability of FISH-Tuning across different pre-trained models, confirming that FISH-Tuning is effective and robust in enhancing the performance of PEFT methods beyond the standard BERT settings.
5.4.4 Contrastive Study
We further investigate the impact of the FISH Mask by conducting a contrastive study on the LoRA framework, as detailed in Table 5 in Appendix B. In this study, we compare the standard LoRA method with three variants: our proposed LoRA-FISH (the standard FISH-Tuning method), a variant where important parameters are selected at random (LoRA-FISH-rand), and another variant where the importance ordering is reversed (LoRA-FISH-rev). The parameter selection method of LoRA-FISH-rev is described as follows, replacing Equation 7:
$\theta_{\text{selected}}=\{\theta_{i}\ |\ \hat{F}_{\theta_{i}}\leq\text{sort\_reverse}(\hat{F}_{\theta})_{k}\}$ (18)
At a low trainable parameter ratio of 0.0057%, LoRA-FISH improves the average score over the original LoRA—from 68.45 to 68.90. In contrast, while the random selection variant (LoRA-FISH-rand) attains a comparable average (68.74), the reverse-ordering strategy (LoRA-FISH-rev) leads to a noticeable drop, yielding an average score of only 66.71. As the parameter ratio increases to 0.0099%, the superiority of the FISH Mask selection becomes more pronounced. LoRA-FISH boosts the average score to 72.25 compared to 70.42 for the original method, whereas LoRA-FISH-rand and LoRA-FISH-rev obtain lower averages of 69.62 and 67.93, respectively.
A similar pattern is also observed at higher ratios. At 0.0142%, LoRA-FISH achieves an average score of 72.64, outperforming both the original LoRA (72.08) and the two contrastive variants (71.89 for the random selection and 71.86 for the reverse ordering). At the highest ratio tested (0.0184%), although both LoRA-FISH and its random counterpart yield comparable averages (72.91 and 72.92, respectively), the reverse ordering variant still falls slightly behind (72.65), and all three methods surpass the original method (72.11).
These findings reinforce that the specific parameter importance guided by the FISH Mask is significant: while randomly or reversed selection strategies may occasionally yield competitive performance at certain scales, they consistently fail to match the robust improvements delivered by FISH Mask approach.
6 Conclusion
In this paper, we propose FISH-Tuning, a new method that uses the selective PEFT method in the Addition-based and Reparameterization-based PEFT methods. More concretely, we integrate the FISH Mask into LoRA, DoRA, Adapter, Prefix-Tuning, and $(\text{IA})^{3}$. With the same ratio of trainable parameters, our method outperforms the original PEFT method most of the time. In future work, we will explore our approach in Computer Vision, Multi-modality, and Quantization.
Limitations
There are still some questions with FISH-Tuning not addressed in this paper:
1) FISH-Tuning does not perform well within the UniPELT framework. We hypothesize that the reason for this phenomenon is that UniPELT uses a gating value, ${\mathcal{G}_{m}}\in(0,1)$. Each individual PEFT module might possess "emergent abilities" similar to causal language models Wei et al. (2022). While an individual PEFT module can perform well using FISH-Tuning, when using UniPELT, the weights of each PEFT module are multiplied by a gating value, ${\mathcal{G}_{m}}\in(0,1)$. This gating value may "dilute" the individual PEFT module’s "emergent abilities". This explanation is merely our hypothesis and requires further experimental validation.
2) We use the LoRA method as an example. To achieve the same ratio of trainable parameters in FISH-Tuning, we can select different trainable layers and LoRA ranks. Tables 3 and 4 in Appendix A demonstrate our experiments on this. According to the results, we have not found any obvious pattern for determining how to select the suitable ranks and layers to achieve the best result.
Nevertheless, we are excited to see the huge potential of FISH-Tuning already demonstrated in existing experiments and look forward to more tests and suggestions from the community.
Ethics Statement
This work aims to contribute to the advancement of Machine Learning. We encourage ethical and responsible application of our results to prevent societal harm. To ensure transparency and reproducibility, and to promote trust and integrity in Machine Learning research, we will release our code and methods publicly.
Acknowledgments
We use generative AI tools only to assist with the language of the paper. We only use them to check for grammatical issues. All new ideas and new text are written by ourselves.
Appendix A LoRA Different Layer
Method Trainable Parameters CoLA MRPC RTE STS-B WNLI Avg
Original-LoRA 0.0057% 43.27 80.94 58.48 84.94 53.52 64.23
LoRA-FISH-rk1-lay5 0.0057% 44.28 80.25 58.48 86.41 53.52 64.59
LoRA-FISH-rk1-lay6 0.0057% 44.19 84.73 58.48 86.93 53.52 65.57
LoRA-FISH-rk1-lay8 0.0057% 46.92 84.66 58.48 87.12 53.52 66.14
LoRA-FISH-rk1-lay10 0.0057% 46.05 80.01 58.48 86.99 53.52 65.01
LoRA-FISH-rk1-lay12 0.0057% 46.99 84.53 58.48 87.11 53.52 66.13
[…]
LoRA-FISH-rk1-lay5 0.0099% 51.21 85.74 66.06 86.84 53.52 68.67
LoRA-FISH-rk1-lay6 0.0099% 51.23 85.43 64.98 87.06 53.52 68.45
LoRA-FISH-rk1-lay8 0.0099% 52.61 85.71 58.48 87.23 53.52 67.51
LoRA-FISH-rk1-lay10 0.0099% 53.41 80.23 65.70 87.37 53.52 68.05
LoRA-FISH-rk1-lay12 0.0099% 51.03 84.99 58.48 87.44 53.52 67.09
[…]
LoRA-FISH-rk1-lay5 0.0142% 53.58 85.56 65.70 86.79 53.52 69.03
LoRA-FISH-rk1-lay6 0.0142% 53.06 87.02 65.70 86.96 53.52 69.25
LoRA-FISH-rk1-lay8 0.0142% 55.21 83.47 64.62 87.23 53.52 68.81
LoRA-FISH-rk1-lay10 0.0142% 56.32 85.74 66.43 87.21 53.52 69.84
LoRA-FISH-rk1-lay12 0.0142% 53.44 83.33 65.70 87.36 53.52 68.67
[…]
LoRA-FISH-rk1-lay5 0.0184% 55.96 84.18 67.51 86.81 53.52 69.60
LoRA-FISH-rk1-lay6 0.0184% 53.31 85.80 66.06 87.02 53.52 69.15
LoRA-FISH-rk1-lay8 0.0184% 53.67 85.63 65.70 87.25 53.52 69.15
LoRA-FISH-rk1-lay10 0.0184% 55.41 86.68 66.06 87.10 53.52 69.75
LoRA-FISH-rk1-lay12 0.0184% 55.91 86.30 66.06 87.34 53.52 69.83
Table 3: Performance of different methods on different datasets. We set the LoRA rank to 1 and select different layers as trainable parameters.
Method Trainable Parameters CoLA MRPC RTE STS-B WNLI Avg
Original-LoRA 0.0057% 43.27 80.94 58.48 84.94 53.52 64.23
LoRA-FISH-rk1-lay5 0.0057% 44.28 80.25 58.48 86.41 53.52 64.59
LoRA-FISH-rk2-lay5 0.0057% 44.98 83.46 62.45 86.18 53.52 66.12
LoRA-FISH-rk4-lay5 0.0057% 44.75 81.95 58.48 86.67 53.52 65.08
LoRA-FISH-rk8-lay5 0.0057% 44.13 81.28 66.06 86.94 53.52 66.39
LoRA-FISH-rk16-lay5 0.0057% 44.41 83.29 63.54 87.41 53.52 66.43
LoRA-FISH-rk32-lay5 0.0057% 48.09 85.31 65.34 87.50 53.52 67.95
[…]
LoRA-FISH-rk1-lay5 0.0099% 51.21 85.74 66.06 86.84 53.52 68.67
LoRA-FISH-rk2-lay5 0.0099% 51.56 82.61 62.45 86.72 53.52 67.37
LoRA-FISH-rk4-lay5 0.0099% 50.76 83.59 64.26 86.54 53.52 67.74
LoRA-FISH-rk8-lay5 0.0099% 49.95 83.79 65.70 87.04 53.52 68.00
LoRA-FISH-rk16-lay5 0.0099% 47.16 85.15 63.90 87.03 53.52 67.35
LoRA-FISH-rk32-lay5 0.0099% 51.01 85.24 64.98 87.61 53.52 68.47
[…]
LoRA-FISH-rk1-lay5 0.0142% 53.58 85.56 65.70 86.79 53.52 69.03
LoRA-FISH-rk2-lay5 0.0142% 52.56 85.70 66.06 86.50 53.52 68.87
LoRA-FISH-rk4-lay5 0.0142% 51.47 83.89 61.73 86.52 53.52 67.43
LoRA-FISH-rk8-lay5 0.0142% 50.00 85.42 64.98 86.93 53.52 68.17
LoRA-FISH-rk16-lay5 0.0142% 52.67 85.24 64.62 87.09 53.52 68.63
LoRA-FISH-rk32-lay5 0.0142% 52.15 84.26 64.98 87.42 53.52 68.47
[…]
LoRA-FISH-rk1-lay5 0.0184% 55.96 84.18 67.51 86.81 53.52 69.60
LoRA-FISH-rk2-lay5 0.0184% 53.21 85.10 68.95 86.65 53.52 69.49
LoRA-FISH-rk4-lay5 0.0184% 52.03 84.24 63.90 86.61 53.52 68.06
LoRA-FISH-rk8-lay5 0.0184% 50.67 84.70 64.98 86.87 53.52 68.15
LoRA-FISH-rk16-lay5 0.0184% 51.92 84.66 70.40 87.18 53.52 69.54
LoRA-FISH-rk32-lay5 0.0184% 53.17 83.91 67.15 87.47 53.52 69.04
Table 4: Performance of different methods on different datasets. We set the trainable layers to 5 and select different LoRA ranks.
We select different trainable layers and LoRA ranks. The results in Tables 3 and 4.
Appendix B LoRA random and LoRA reverse selection
We compare the standard LoRA method with three variants: our proposed LoRA-FISH (the standard FISH-Tuning method), a variant where important parameters are selected at random(LoRA-FISH-rand), and another variant where the importance ordering is reversed (LoRA-FISH-rev). The results in Tables 5.
Method Trainable Parameters CoLA MRPC RTE SST-2 STS-B WNLI Avg
Original-LoRA 0.0057% 43.27 80.94 58.48 89.56 84.94 53.52 68.45
LoRA-FISH 0.0057% 44.28 80.25 58.48 90.48 86.41 53.52 68.90
LoRA-FISH-rand 0.0057% 46.18 78.38 58.12 89.56 86.69 53.52 68.74
LoRA-FISH-rev 0.0057% 38.91 74.80 58.12 88.99 85.91 53.52 66.71
[…]
LoRA-FISH 0.0099% 51.21 85.74 66.06 90.14 86.84 53.52 72.25
LoRA-FISH-rand 0.0099% 48.67 81.25 58.12 89.22 86.92 53.52 69.62
LoRA-FISH-rev 0.0099% 44.53 74.80 58.12 89.91 86.68 53.52 67.93
[…]
LoRA-FISH 0.0142% 53.58 85.56 65.70 90.71 86.79 53.52 72.64
LoRA-FISH-rand 0.0142% 52.07 82.24 66.43 90.14 86.95 53.52 71.89
LoRA-FISH-rev 0.0142% 48.68 84.03 66.43 91.17 87.33 53.52 71.86
[…]
LoRA-FISH 0.0184% 55.96 84.18 67.51 89.45 86.81 53.52 72.91
LoRA-FISH-rand 0.0184% 54.59 85.49 66.79 89.91 87.23 53.52 72.92
LoRA-FISH-rev 0.0184% 54.10 83.31 67.51 90.25 87.20 53.52 72.65
Table 5: Performance of different methods on different datasets. In each dashed-line area, the first row represents the original method, the second row represents our method, the third row represents the method where we randomly select the important parameters without using the FISH Mask, and the fourth row represents the method where we select the important parameters in reverse order compared to the second method.
Appendix C Hyperparameter settings of the experiment
The hyperparameter settings in our experiment. The result in Table 6.
Hyperparameters value
batch size 32
learning rate 5e-5, except WNLI 5e-6
epoch 400
optimizer Adam
early stop Yes
seed 42
warm up ratio 0.0
number of FISH Mask samples 128
dataset used for Fisher estimation train set
Prefix-Tuning’s prefix length $l$ 30
LoRA $\alpha$ 2 * rank
LoRA dropout rate 0.0
Selected LoRA weights For BERT, we select $W_{Q}$, $W_{K}$, and $W_{V}$.
For ModernBERT and LLaMA-3.2-1B, we select $W_{O}$.
max sequence length details in Table 7
Table 6: The hyperparameters in our experiment.
Task max_seq_length
SST-2 128
CoLA 128
RTE 384
STS-B 256
WNLI 128
MRPC 128
Table 7: The different max_seq_length in different tasks.